A loan of ​$930 was repaid at the end of 18 months with a check for ​$950. What annual rate of interest was​ charged?
Answer by math_tutor2020(3817)


$930 was borrowed, and $950 was repaid
interest = 950-930 = 20 dollars

18 months = (18/12) years = 1.5 years

If we're dealing with simple interest, then,
i = P*r*t
20 = 930*r*1.5
20 = 1395r
r = 20/1395
r = 0.0143 approximately

Answer: Approximately 1.43%